OpenStudy (liv1234):

I need help with a question! Please.(:

OpenStudy (liv1234):

@haleyelizabeth2017

OpenStudy (haleyelizabeth2017):

f(x)=(x-3)^2-3 Right you do (x-_), left you do (x+_) up you do (x)^2+_, down you do (x)^2-_

OpenStudy (liv1234):

Is that all I have to do? Or is there more?

OpenStudy (haleyelizabeth2017):

I believe that is all.

OpenStudy (haleyelizabeth2017):

Wait

OpenStudy (liv1234):

Oh wow, thank you very much!(:

OpenStudy (liv1234):

Okay.

OpenStudy (haleyelizabeth2017):

It's g(x) for the one I did...not f(x) XD so g(x)=(x-3)^2-3

OpenStudy (liv1234):

So, write g(x) instead?

OpenStudy (liv1234):

And that's all I have to write?

OpenStudy (haleyelizabeth2017):

Yeah

OpenStudy (haleyelizabeth2017):

well, the g(x)=______ function XD

OpenStudy (liv1234):

What am I supposed to write in the blank space? cx

OpenStudy (liv1234):

I'm confused, so after I write what you told me before, I put "g(x) = ____ function", but what goes in the space?

OpenStudy (haleyelizabeth2017):

g(x)=(x-3)^2-3